Evaluate the integral ∫¹ₒ [x²⁄(√4-x²)ᶟ]dx
1
Expert's answer
2020-06-29T19:05:27-0400

To solve this integration, we use the method of integration " Trigonometric Substitution "

Let "x=2sin \theta" , then "dx=2cos \theta d\theta"

Note that when "x=0" then "\theta =0"

and when "x=1" , then "\theta=\pi/6"


Let "I=\int\limits_{0}^{1} \frac { x^{2}} {(\sqrt4-x^{2} )^{3}} dx" , then "I=\int\limits_{0}^{\pi/6} \frac { 4sin^{2}x} {8cos^{3}\theta} *2cos \theta d\theta =\int\limits_{0}^{\pi/6} tan^{2}\theta d\theta"


Thus "I=\int\limits_{0}^{\pi/6} (sec^{2}\theta-1)d\theta=[tan\theta-\theta]^{\pi/6}_{0}"

and then "I=[tan\theta-\theta]^{\pi/6}_{0}=(1/\sqrt3)-\pi/6"


Therefore we have that

"\int\limits_{0}^{1} \frac { x^{2}} {(\sqrt4-x^{2} )^{3}} dx=(1/\sqrt3)-(\pi/6)"